OpenStudy (anonymous):

HELP AND PLEASE WORK OUT SO I UNDERSTAND HOW TO DO IT. 2x+7y=1 x+5y=2

OpenStudy (amistre64):

any initial thoughts on how to approach it?

OpenStudy (amistre64):

if we take the 2nd equation, we can multiply it by 2 to see what 2x is equivalent to

OpenStudy (amistre64):

2(x+5y=2) 2x+10y = 4 - 10y -10y ------------- 2x = 4 - 10y everywhere we see a 2x, we can replace it by 4-10y then see if there is a way to find a solution for y

OpenStudy (anonymous):

2x=4-10y? what next what does x = ?

OpenStudy (amistre64):

forget about x for the moment, and replace all the 2x parts you see with 4-10y that should give you an idea on what to do next

OpenStudy (anonymous):

OK. 2x=4-10y is the problem finished?

OpenStudy (amistre64):

of course not, there is more to do which is why i suggested that you actually do more. 2x + 7y = 1 ^^ replace it 4 - 10y + 7y = 1

OpenStudy (amistre64):

do you agree that this is solvable for y now?

OpenStudy (anonymous):

\[x +5y=2\] \[x=2-5y\] then substitute the x to 2x+7y=1 \[2(2-5y)+7y=1\] \[4-10y+7y=1\] \[4-3y=1\]

OpenStudy (anonymous):

\[3y=4-1\] \[3y=3\] \[y=1\] now you have the value of y

OpenStudy (anonymous):

so you can substitute it to x+5y=2 to find the x

OpenStudy (anonymous):

ok this is what i have on my paper. 2x+7y=1 (x+5y=2)2 now I'm really confused.

OpenStudy (amistre64):

multiply the 2 thru the paranthesis ... distribute it 2x+7y=1 (x+5y=2)2 2x+7y=1 2x+10y=4 now there are a few ways to approach this, substitution or elimination ... both work out fine

OpenStudy (anonymous):

well i don't know what to do after distributive

OpenStudy (amistre64):

well, my thought was to solve for 2x, and run a substitution, which i worked thru already. tblue even worked thru a similar substitution process and determined a specific y value for you.

OpenStudy (amistre64):

the key strategy is simply getting rid of 2 variables so that you only have 1 variable to solve .... and a 1 variable solution is gone over extensively prior to these systems. 4 - 10y + 7y = 1 we can use that to solve for a specific value of y

OpenStudy (amistre64):

if we use the elimination method, we are primed for getting rid of the x parts by subtracting one equation from the other 2x+7y=1 -(2x+10y=4) 2x +7y = 1 -2x -10y = -4 -------------- 0 -3y = -3 -3y = -3 when y = ?

OpenStudy (amistre64):

good luck with this, i have to be going